OpenStudy (anonymous):

how would you solve this? tan(theta) + 3cot(theta) = 5sec(theta)

ganeshie8 (ganeshie8):

may be change everything into sin/cos

ganeshie8 (ganeshie8):

and see if u can get a quadratic/cubic in sin or cos alone

OpenStudy (anonymous):

good luck i checked wolfram and \(\frac{\pi}{6}\) works, but i am not sure how to get it other than checking it

OpenStudy (anonymous):

i have gotten up to here 2*sin^2(theta) + 5*sin(theta) - 3 = 0

ganeshie8 (ganeshie8):

i see a quadratic in sin... dive in and give it a try

ganeshie8 (ganeshie8):

that looks right ^

OpenStudy (anonymous):

oh that looks promising but there are others that you might miss

ganeshie8 (ganeshie8):

use quadratic formula or factor it

ganeshie8 (ganeshie8):

2*sin^2(theta) + 5*sin(theta) - 3 = 0 analogous to : 2x^2 + 5x - 3 = 0

OpenStudy (anonymous):

which, by some miracle, factors!